### v1

@@ 4.2. Helmholtz-Weyl decomposition of $\mathbf{W}^{m,p}_{t}(\Gamma)$
guaranteed by 3.16. Furthermore, $\|\uppi\|_{1,p;\Gamma}\leq C\|\mathbf{v}\|_{0,p;\Gamma}$.
@@
$\|\mathbf{h}\|_{0,p;\Gamma}+\|\nabla_{\Gamma}\uppi\|_{0,p;\Gamma}\leq\|\mathbf{v}\|_{0,p;\Gamma}+\|\mathbf{g}\|_{0,p;\Gamma}+\|\nabla_{\Gamma}\uppi\|_{0,p;\Gamma}\leq C\|\mathbf{v}\|_{0,p;\Gamma}.$
To prove (4.3) for $m\geq 1$, it is enough to notice that if $\mathbf{v}\in\mathbf{W}^{m,p}_{t}(\Gamma)$, then the right-hand side $\operatorname{div}_{\Gamma}\mathbf{v}$ of (4.4) lies in $\mathrm{W}^{m-1,p}_{\#}(\Gamma)$, and hence by the $\mathrm{L}^{p}$-based regularity of the Laplace–Beltrami operator $\Delta_{\Gamma}$ (cf. Lemmas 3.1, 3.2 and 3.3) we know that $\uppi\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and $\|\uppi\|_{m+1,p;\Gamma}\leq C\|\operatorname{div}_{\Gamma}\mathbf{v}\|_{\mathrm{W}^{m-1,p}(\Gamma)}\leq C\|\mathbf{v}\|_{m,p,\Gamma}$. Consequently, $\mathbf{g}=\nabla_{\Gamma}\uppi\in\nabla_{\Gamma}\mathrm{W}^{m+1,p}_{\#}(\Gamma)\subseteq\mathbf{W}^{m,p}_{t}(\Gamma)$ and so $\mathbf{h}=\mathbf{v}-\mathbf{g}\in\mathbf{W}^{m,p}_{t}(\Gamma)$. That is, $\mathbf{v}=\mathbf{h}+\mathbf{g}$ with the corresponding stability bound.
The uniqueness of the decomposition is proved as follows. Suppose $\mathbf{v}\in\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)\cap\nabla_{\Gamma}\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Then, $\mathbf{v}=\nabla_{\Gamma}\phi$ for some $\phi\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and $\operatorname{div}_{\Gamma}\mathbf{v}=\Delta_{\Gamma}\phi=0$. Hence, by 3.16 (well-posedness of the Laplace–Beltrami in $\mathrm{W}^{1,p}_{\#}(\Gamma)$), $\phi=0$ and so $\mathbf{v}=\textbf{0}$.
@@
Definition 4.4 (Leray projection on $\mathbf{L}^{p}_{t}(\Gamma)$).
Let $1<p<\infty$. We define the Leray projection $P_{\sigma}:\mathbf{L}^{p}_{t}(\Gamma)\rightarrow\mathbf{L}^{p}_{t,\sigma}(\Gamma)$ by $P_{\sigma}\mathbf{v}=\mathbf{h}$ for each $\mathbf{v}=\mathbf{h}+\mathbf{g}\in\mathbf{L}^{p}_{t}(\Gamma)$, where $\mathbf{h}\in\mathbf{L}^{p}_{t,\sigma}(\Gamma)$ and $\mathbf{g}\in\nabla_{\Gamma}\mathrm{W}_{\#}^{1,p}(\Gamma)$ are the unique components of $\mathbf{v}$ given by 4.3 (Helmholtz-Weyl decomposition on $\Gamma$). It shall also prove useful to denote $P_{\sigma}^{\perp}\mathbf{v}=\mathbf{g}$.
@@
$\|P_{\sigma}\mathbf{v}\|_{m,p,\Gamma}+\|P_{\sigma}^{\perp}\mathbf{v}\|_{m,p,\Gamma}\leq C\|\mathbf{v}\|_{m,p,\Gamma},\qquad\mathbf{v}\in\mathbf{W}^{m,p}_{t}(\Gamma).$ (4.6)
@@
$\int_{\Gamma}\mathbf{v}\cdot\mathbf{w}=\int_{\Gamma}P_{\sigma}\mathbf{v}\cdot P_{\sigma}\mathbf{w}+\int_{\Gamma}P_{\sigma}^{\perp}\mathbf{v}\cdot P_{\sigma}^{\perp}\mathbf{w}.$ (4.7)
@@ 4.3. Well-posedness in $\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$
In this section we establish the well-posedness of a variational formulation for the tangent Stokes equations (S) in a non-Hilbert setting. Indeed, let $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. For each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$ consider the problem: find $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
Problem (4.8) is of the generalized Babuška–Brezzi-type (see A.3). Its well-posedness hinges on a collection of inf-sup conditions, which we prove below. Strikingly, due to 2.1, we can prove the well-posedness of (4.8) for all $p>1$ at once using the well-posedness and higher-regularity results for the Laplace–Beltrami and Bochner–Laplace problems developed earlier.
The following two results establish the bijectivity of $\Delta_{B}:=-\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ regarded as an operator acting from $\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ to $(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ (see A.2).
Proposition 4.5 (injectivity of the associated adjoint operator).
Let $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{3}$. If $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$ satisfies
@@
$\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{u}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v},\qquad\forall\,\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma).$ (4.10)
By virtue of 4.3 (Helmholtz decomposition), each $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ can be uniquely decomposed as $\mathbf{u}=\mathbf{w}+\nabla_{\Gamma}\phi$, where $\mathbf{w}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ and $\phi\in\mathrm{W}^{2,p}_{\#}(\Gamma)$. Consequently, and employing the assumption (4.9) on $\mathbf{v}$ for $\mathbf{u}\leftarrow\mathbf{w}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$, problem (4.10) is equivalent to finding $q\in\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ such that
@@
Problem (4.11) is an instance of an ultraweak formulation (3.11) for the Laplace–Beltrami problem, whose well-posedness follows from the continuous bijectivity of the Laplace–Beltrami operator from $\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ onto $\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ (cf. Lemmas 3.1, 3.2 and 3.3). Hence, there is a unique $q\in\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ that satisfies (4.11) (and hence also (4.10)). If $p^{*}\geq 2$, then 4.2 ($\mathrm{L}^{2}$-based well-posedness of the tangent Stokes problem) applied to the context given by (4.10) and $\operatorname{div}_{\Gamma}\mathbf{v}=0$, yields $\mathbf{v}=\textbf{0}$.
Otherwise if $1<p^{*}<2$, based on Sobolev embeddings (cf. Section 2.2), we proceed by a bootstrapping argument to increase the Sobolev-integrability parameter of the pair $(\mathbf{v},q)$. First, by utilizing 2.1 (valid by density) and the fact that $\operatorname{div}_{\Gamma}\mathbf{v}=0$, we can rewrite (4.11) as
@@
where $h[\mathbf{v}]:=\operatorname{div}_{\Gamma}\left((\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}\right)\in\mathrm{L}^{p^{*}}(\Gamma)$. Therefore, we deduce from the $\mathrm{L}^{p}$-based ($1<p<\infty$) Sobolev regularity theory of the Laplace–Beltrami operator (see Section 3) that $q\in\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ is in fact in $\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$. With this further regularity at hand, we now go back to (4.10) and integrate by parts on the left-hand side to obtain
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}=-\int_{\Gamma}\nabla_{\Gamma}q\cdot\mathbf{u},\qquad\forall\,\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma).$ (4.13)
Hence, since $\nabla_{\Gamma}q\in\mathrm{L}^{p^{*}}_{t}(\Gamma)$, we deduce from the $\mathrm{L}^{p}$-based ($1<p<\infty$) Sobolev regularity theory of the Bochner–Laplace operator (see Section 3) that $\mathbf{v}\in\mathbf{W}^{2,p^{*}}_{t}(\Gamma)$. More precisely, since $p^{*}<2\leq d$, from 2.6 (Gagliardo–Nirenber–Sobolev inequality) we deduce that $h[\mathbf{v}]\in\mathrm{L}^{p_{0}^{*}}(\Gamma)$ for $p_{0}^{*}:=\frac{dp^{*}}{d-p^{*}}$. Consequently, the Sobolev regularity of the Laplace–Beltrami and Bochner–Laplace operator applied to (4.12) and (4.13) yields $\phi\in\mathrm{W}^{2,p_{0}^{*}}_{\#}(\Gamma)$ and $\mathbf{v}\in\mathbf{W}^{2,p_{0}^{*}}_{t}(\Gamma)$.
We can iterate this procedure as follows: if $p_{n}^{*}<2$, define
$p_{n+1}^{*}:=\frac{dp_{n}^{*}}{d-p_{n}^{*}}.$ (4.14)
Clearly, the sequence $\{p_{n}^{*}\}_{n\in\mathrm{N}}$ is nondecreasing. We claim that there exists $N\in\mathrm{N}$ such that $p_{N}^{*}\geq 2$. In fact, if we assume otherwise, then the monotonicity of $p_{n}^{*}$ yields the existence of $L\geq 1$ such that $\lim_{n\to\infty}p_{n}^{*}=L$. By taking limit as $n\to\infty$ in (4.14), we obtain that $L:=\frac{dL}{d-L}$, or equivalently, $L=0$; which is a contradiction. Hence, $(\mathbf{v},\phi)\in\mathbf{W}^{2,p_{N}^{*}}_{t}(\Gamma)\times\mathrm{W}^{1,p_{N}^{*}}_{\#}(\Gamma)$, and so by 4.2, $\mathbf{v}=\textbf{0}$. ∎
Proposition 4.6 (surjectivity of the associated adjoint operator).
Let $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{3}$. Then, there exists a positive constant $\alpha>0$ such that
@@
Similarly to the proof of 3.13 and 3.19, it suffices to show
@@
from which (4.15) follows by contradiction, using compactness arguments (Banach–Alaoglu Theorem and compact Sobolev embeddings) and 4.5. We omit further details and simply proceed to show (4.16).
By using properties of the Leray projection (cf. 4.4) we write in first instance,
@@
Recalling that $P_{\sigma}^{\perp}\mathbf{v}\in\nabla_{\Gamma}\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ for each $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$, we have by integrating by parts and 2.1 with $\nabla_{\Gamma}\phi\leftarrow P_{\sigma}^{\perp}\mathbf{v}$ and $\mathbf{v}\leftarrow\mathbf{u}$ (valid by density) that
@@
for each $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$. By utilizing this last identity, the $\mathbf{W}^{1,p^{*}}(\Gamma)$-stability of $P_{\sigma}^{\perp}$ and the inf-sup condition for the Bochner-Laplacian 3.19, we arrive at
@@
This finishes the proof. ∎
The following inf-sup condition encodes, for arbitrary $1<p<\infty$, the surjectivity of the covariant divergence $\operatorname{div}_{\Gamma}$ as an operator acting from $\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$ onto $(\mathrm{L}^{p}_{\#}(\Gamma))^{\prime}\equiv\mathrm{L}^{p^{*}}_{\#}(\Gamma)$.
Let $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{2}$. Then, there is a constant $c>0$ depending on $\Gamma$ and $p$ such that
Proof.
The proof is a generalization of the proof of 4.1 for a general $\mathrm{L}^{p}$-based setting.
Fix $\uppi\in\mathrm{L}_{\#}^{p}(\Gamma)$. By the $\mathrm{L}^{p}$-based regularity theory of the Laplace–Beltrami operator (cf. 3.1, 3.2, 3.3) there exists a unique $\phi\in\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ such that $\Delta_{\Gamma}\phi=|\uppi|^{p-2}\uppi-\alpha_{\uppi}$ on $\Gamma$, where $\alpha_{\pi}:=\frac{1}{|\Gamma|}\int_{\Gamma}|\uppi|^{p-2}\uppi$. Furthermore, $\|\phi\|_{\mathrm{W}^{2,p^{*}}(\Gamma)}\lesssim\||\uppi|^{p-2}\uppi-\alpha_{\uppi}\|_{0,p^{*};\Gamma}\lesssim\|\uppi\|_{0,p;\Gamma}^{p-1}$. We use the function $\mathbf{w}:=\nabla_{\Gamma}\phi\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$, that satisfies $\operatorname{div}_{\Gamma}\mathbf{w}=|\uppi|^{p-2}\uppi-\alpha_{\uppi}$ on $\Gamma$ and $\|\mathbf{w}\|_{1,p^{*};\Gamma}\lesssim\|\phi\|_{\mathrm{W}^{2,p^{*}}(\Gamma)}\lesssim\|\uppi\|_{0,p;\Gamma}^{p-1}$, to bound the supremum from below. This yields
$\sup_{\begin{subarray}{c}\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)\\ \mathbf{v}\neq\mathbf{0}\end{subarray}}\frac{\left|\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}\right|}{\|\mathbf{v}\|_{1,p^{*};\Gamma}}\geq\frac{\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{w}}{\|\mathbf{w}\|_{1,p^{*};\Gamma}}\gtrsim\frac{\|\uppi\|_{0,p;\Gamma}^{p}-\alpha_{\uppi}\int_{\Gamma}\pi}{\|\uppi\|_{0,p;\Gamma}^{p-1}}=\|\uppi\|_{0,p;\Gamma},$ (4.19)
which finishes the proof. ∎
@@
Let $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Assume $\Gamma$ is of class $C^{3}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$ there exist a unique $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ solution of (4.8). Moreover, the exists a constant $C>0$, depending only on $\Gamma$ and $p$, such that
@@
We first notice that all bilinear forms defining problem (4.8) are bounded in their respective domains, and that for each $1<r<\infty$, the set $V:=\left\{\mathbf{w}\in\mathbf{W}^{1,r}_{t}(\Gamma):\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{w}=0,\quad\forall q\in\mathrm{L}^{r^{*}}_{\#}(\Gamma)\right\}$ reduces to $V=\mathbf{W}^{1,r}_{t,\sigma}(\Gamma)$. Then, thanks to 4.5, 4.6 and 4.7, the proof follows from a straightforward application of A.3 (generalized Babuška–Brezzi theory in reflexive Banach spaces). ∎

### v2

@@ 4.1. Helmholtz–Weyl decomposition of $\mathbf{W}^{m,p}_{t}(\Gamma)$
guaranteed by 2.15. We know that $\|\uppi\|_{1,p;\Gamma}\lesssim\|\mathbf{v}\|_{0,p;\Gamma}$.
@@
$\|\mathbf{h}\|_{0,p;\Gamma}+\|\nabla_{\Gamma}\uppi\|_{0,p;\Gamma}\leq\|\mathbf{v}\|_{0,p;\Gamma}+2\|\nabla_{\Gamma}\uppi\|_{0,p;\Gamma}\lesssim\|\mathbf{v}\|_{0,p;\Gamma}.$
The case $m\geq 1$ follows by integrating the right-hand side of (4.5) by parts and using the higher regularity result within 2.15. We omit further details.
For any $m\geq 0$, the uniqueness of the decomposition is proved as follows. Suppose $\mathbf{v}\in\mathbf{W}^{m,p}_{t,\sigma}(\Gamma)\cap\nabla_{\Gamma}\mathrm{W}^{m+1,p}_{\#}(\Gamma)$. Then, $\mathbf{v}=\nabla_{\Gamma}\phi$ for some $\phi\in\mathrm{W}^{m+1,p}_{\#}(\Gamma)$ and $\operatorname{div}_{\Gamma}\mathbf{v}=\Delta_{\Gamma}\phi=0\in(\mathrm{W}^{1,p^{*}}(\Gamma))^{\prime}$; that is, $\int_{\Gamma}\nabla_{\Gamma}\phi\cdot\nabla_{\Gamma}q=0$ for each $q\in\mathrm{W}^{1,p^{*}}(\Gamma)$. Hence, by 2.15, $\phi=0$ and so $\mathbf{v}=\textbf{0}$.
@@
Definition 4.4 (Leray projection in $\mathbf{L}^{p}_{t}(\Gamma)$).
Under the hypotheses of 4.3, we define the Leray projection $P_{\sigma}:\mathbf{L}^{p}_{t}(\Gamma)\rightarrow\mathbf{L}^{p}_{t,\sigma}(\Gamma)$ by $P_{\sigma}\mathbf{v}=\mathbf{h}$ for each $\mathbf{v}=\mathbf{h}+\mathbf{g}\in\mathbf{L}^{p}_{t}(\Gamma)$, where $\mathbf{h}\in\mathbf{L}^{p}_{t,\sigma}(\Gamma)$ and $\mathbf{g}\in\nabla_{\Gamma}\mathrm{W}_{\#}^{1,p}(\Gamma)$ are the unique components of $\mathbf{v}$ given by 4.3 (Helmholtz–Weyl decomposition on $\Gamma$). It shall also prove useful to denote $P_{\sigma}^{\perp}\mathbf{v}=\mathbf{g}$.
@@
$\|P_{\sigma}\mathbf{v}\|_{m,p;\Gamma}+\|P_{\sigma}^{\perp}\mathbf{v}\|_{m,p;\Gamma}\leq C\|\mathbf{v}\|_{m,p;\Gamma},\qquad\mathbf{v}\in\mathbf{W}^{m,p}_{t}(\Gamma).$ (4.6)
@@
$\int_{\Gamma}\mathbf{v}\cdot\mathbf{w}=\int_{\Gamma}P_{\sigma}\mathbf{v}\cdot P_{\sigma}\mathbf{w}+\int_{\Gamma}P_{\sigma}^{\perp}\mathbf{v}\cdot P_{\sigma}^{\perp}\mathbf{w}.$
@@ 4.2. Well-posedness in $\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$
In this section we verify that problem (4.1) verifies the conditions of A.1 (generalized Babuška–Brezzi). The following two results establish the bijectivity of $\Delta_{B}:=-\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ regarded as an operator acting from $\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ to $(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ (see [5, Theorem A.2]).
Proposition 4.5 (injectivity of the adjoint operator).
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. If $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$ satisfies
@@
$\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{u}=\int_{\Gamma}\nabla_{\Gamma}\mathbf{v}:\nabla_{\Gamma}\mathbf{u},\qquad\forall\,\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma).$ (4.8)
By virtue of 4.3 (Helmholtz–Weyl decomposition on $\Gamma$), each $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ in (4.8) can be uniquely decomposed as $\mathbf{u}=\mathbf{w}+\nabla_{\Gamma}\phi$, where $\mathbf{w}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ and $\phi\in\mathrm{W}^{2,p}_{\#}(\Gamma)$. Consequently, employing the assumption (4.7) for $\mathbf{u}$ replaced by $\mathbf{w}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$, problem (4.8) is equivalent to finding $q\in\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ such that
@@
Problem (4.9) is an instance of an ultraweak formulation (2.10) for the Laplace–Beltrami problem, whose well-posedness is ensured by 2.17 (ultra-weak solutions of the Laplace–Beltrami operator on $\Gamma$). Therefore, there is a unique $q\in\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ that satisfies (4.9), whence also (4.8). If $p^{*}\in[2,\infty)$, then 4.2 (well-posedness of tangent Stokes in $\mathbf{H}^{1}_{t}(\Gamma)\times\mathrm{L}^{2}_{\#}(\Gamma)$) applied to the context given by (4.8) and $\operatorname{div}_{\Gamma}\mathbf{v}=0$, yields $\mathbf{v}=\textbf{0}$.
Otherwise if $p^{*}\in(1,2)$, we proceed by a bootstrapping argument based on Sobolev embeddings (cf. Section 2.1) to increase the Sobolev integrability parameter of the pair $(\mathbf{v},q)$. First, by utilizing 2.13 (weak commutator identity), which is valid by density for $\phi\in\mathrm{W}^{2,p}_{\#}(\Gamma)$, integration-by-parts formula (2.5) and the fact that $\operatorname{div}_{\Gamma}\mathbf{v}=0$, we can rewrite (4.9) as
@@
where $h[\mathbf{v}]:=\operatorname{div}_{\Gamma}\left((\operatorname*{tr}(\mathbf{B})\mathbf{B}-\mathbf{B}^{2})\mathbf{v}\right)\in\mathrm{L}^{p^{*}}(\Gamma)$. Therefore, we deduce from the $\mathrm{L}^{p}$-based Sobolev regularity theory of the Laplace–Beltrami operator for $p\in(1,\infty)$ (see 2.17 and 2.15) that $q\in\mathrm{L}^{p^{*}}_{\#}(\Gamma)$ is in fact in $\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$. From 2.8 (Gagliardo–Nirenberg–Sobolev inequality), we infer that $q\in\mathrm{W}^{1,p^{*}_{0}}_{\#}(\Gamma)$, where $p_{0}^{*}:=\frac{dp^{*}}{d-p^{*}}>p^{*}$. With this, we can go back to (4.8) and use 3.6 (well-posedness of the Bochner Laplacian in $\mathbf{W}^{1,p}_{t}(\Gamma)$) to deduce that $\mathbf{v}\in\mathbf{W}^{1,p^{*}_{0}}_{t}(\Gamma)$. Using the improved integrability of $\mathbf{v}$, we consequently obtain the improved integrability for $q\in\mathrm{W}^{2,p^{*}_{0}}_{\#}(\Gamma)$ from (4.10). Iterating this process eventually yields $(\mathbf{v},q)\in\mathbf{H}^{1}_{t}(\Gamma)\times\mathrm{H}^{2}_{\#}(\Gamma)$, which brings us back to the first case. This ends the proof. ∎
Proposition 4.6 (surjectivity of the adjoint operator).
Let $p\in(1,\infty)$ and assume that $\Gamma$ is of class $C^{2}$. Then, there exists a positive constant $\alpha>0$ such that
@@
Step 1: We first establish the (in principle) weaker estimate
@@
which will be later used in Step 2.
Indeed, by using properties of the Leray projection (cf. 4.4) we write in first instance,
@@
Recalling that $P_{\sigma}^{\perp}\mathbf{v}\in\nabla_{\Gamma}\mathrm{W}^{2,p^{*}}_{\#}(\Gamma)$ for each $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$, and applying 2.13 (weak commutator identity) with $\nabla_{\Gamma}\phi\leftarrow P_{\sigma}^{\perp}\mathbf{v}$ and $\mathbf{v}\leftarrow\mathbf{u}$, which is valid by density, yields
@@
for each $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$. By utilizing this last identity, the $\mathbf{W}^{1,p^{*}}(\Gamma)$-stability of $P_{\sigma}^{\perp}$ (cf. (4.6)) and the corresponding inf-sup condition for the Bochner Laplacian given by 3.6 (well-posedness of the Bochner Laplacian in $\mathbf{W}^{1,p}_{t}(\Gamma)$) and [5, Theorem A.1] (characterization of injectivity and surjectivity), we arrive at
@@
This establishes (4.12) and finishes the proof of Step 1.
Step 2: With (4.12) at hand, we now proceed to prove (4.11) by contradiction in the spirit of Peetre–Tartar’s Lemma [26, Lemma A.38].
In fact, assume there exists a sequence $\{\mathbf{u}_{n}\}_{n\in\mathrm{N}}\subseteq\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ such that $\|\mathbf{u}_{n}\|_{1,p;\Gamma}=1$ for each $n\in\mathrm{N}$ and
$\sup_{0\neq\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)}\frac{\left|\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{n}:\nabla_{\Gamma}\mathbf{v}\right|}{\|\mathbf{v}\|_{1,p^{*};\Gamma}}\xrightarrow[]{n\to\infty}0.$
In particular,
$\lim_{n\to\infty}\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}_{n}:\nabla_{\Gamma}\mathbf{v}=0,\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$ (4.13)
Since $\{\mathbf{u}_{n}\}_{n\in\mathrm{N}}$ is bounded in $\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ (which is reflexive), the Banach–Alaoglu Theorem [12, Theorem 3.18] and 2.12 (compact Sobolev embedding) implies the existence of $\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ such that (up to a subsequence)
$\mathbf{u}_{n}\rightharpoonup\mathbf{u}\quad\text{in $\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$},\qquad\mathbf{u}_{n}\to\mathbf{u}\quad\text{in $\mathbf{L}^{p}_{t,\sigma}(\Gamma)$}.$
Thus, for each $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$,
[…]
By 4.5 (injectivity of the adjoint operator), we deduce that $\mathbf{u}=0$ a.e. on $\Gamma$. Hence, passing to the limit $n\to\infty$ in inequality (4.12) for $\mathbf{u}=\mathbf{u}_{n}$ yields $\lim_{n\to\infty}\|\mathbf{u}_{n}\|_{1,p;\Gamma}=0$, which contradicts the fact that $\|\mathbf{u}_{n}\|_{1,p;\Gamma}=1$ for each $n\in\mathrm{N}$. This concludes the proof. ∎
With 4.5 and 4.6 at hand, we now establish the well-posedness of problem (4.1).
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Then, for each $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ and $g\in\mathrm{L}^{p}_{\#}(\Gamma)$ there exist a unique $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ solution of (4.1). Moreover, there exists a constant $C>0$, depending only on $\Gamma$ and $p$, such that
@@
We first notice that all bilinear forms defining problem (4.1) are bounded in their respective domains, and that for each $r\in(1,\infty)$, the set $V:=\left\{\mathbf{w}\in\mathbf{W}^{1,r}_{t}(\Gamma):\int_{\Gamma}q\operatorname{div}_{\Gamma}\mathbf{w}=0,\quad\forall q\in\mathrm{L}^{r^{*}}_{\#}(\Gamma)\right\}$ reduces to $V=\mathbf{W}^{1,r}_{t,\sigma}(\Gamma)$. Then, thanks to 4.5, 4.6 and 4.1, the proof follows from a straightforward application of A.1 (generalized Babuška–Brezzi theory in reflexive Banach spaces). ∎